# Seeding neural network quantum states with tensor network states

### v1

@@ III.2 VMC simulation under open boundary conditions
Figure 11: (a) Energy error as a function of the rank $R$ of the CP decomposition for $h/J=1$ and $n=16$. Filled circles represent the energy error ($|E_{\mathrm{initial,open}}-E_{\mathrm{exact,open}}|/|E_{\mathrm{exact,open}}|$) of the initial state for open boundary conditions used for the VMC method. Up-pointing and down-pointing triangles represent the energy error ($|E_{\mathrm{VMC,open}}-E_{\mathrm{exact,open}}|/|E_{\mathrm{exact,open}}|$, $|E_{\mathrm{VMC,periodic}}-E_{\mathrm{exact,periodic}}|/|E_{\mathrm{exact,periodic}}|$) of the optimized state for open and periodic boundary conditions, respectively. (b) Same as (a) but they are shown in the logarithmic scale. We plot the line proportional to $R^{-2}$ as a reference.
@@ III.3 VMC simulation under periodic boundary conditions
In general, it is hard to prepare the MPS representation of the ground-state wave function under periodic boundary conditions. However, the ground-state wave function should be insensitive to boundary conditions when the system is large enough. Therefore, one may seed the wave function for periodic boundary conditions with the initial wave function for open boundary conditions. We will demonstrate that the initial state obtained by the CP decomposition for the open-boundary system can also be efficiently used for the simulation under periodic boundary conditions.
@@ III.4 Size dependence of the required rank of the CP decomposition
Finally, we numerically investigate how large the rank $R$ of the CP decomposition should be for arbitrary system sizes. We focus on a rather difficult case at the critical point ($h/J=1$) of the transverse-field Ising model under periodic boundary conditions. The ground state is known to exhibit a logarithmic correction in the entanglement entropy, which naturally requires slightly larger bond dimensions in the MPS representation compared to bond dimensions required for conventional gapped ground states. As we will see below, for all systems of $n$ sites, the condition $R=\mathcal{O}(n)$ is sufficient to obtain the ground-state wave function having a small energy error that is independent of the system size.
We first prepare the MPS representation of the ground-state wave function with the maximum bond dimension $D_{\mathrm{max}}=n/2$ for all systems of $n$ sites using the DMRG method. The energy error from the true ground state is found to be less than $10^{-9}$ in the unit of $J$ for any system size that we consider. We then CP decompose the MPS representation using the ALS method with $100$ iterations from the random initial CP factors. We optimize the parameters in the RBM wave function generated by the CP decomposed tensor using the VMC method. We keep the rank $R$ of the CP decomposition to $R=D_{\mathrm{max}}(=n/2)$ for all systems of $n$ sites. The number of variational parameters ($b_{l,k}$ and $W_{l,k}$) is $2nR=n^{2}$, which is much smaller than the dimension of the Hilbert space. We use $10^{6}$ ($4\times 10^{6}$) Monte Carlo steps for evaluating physical quantities during the optimization (during the statistical processing).
@@ IV Summary and outlook
We proposed a method for approximately converting the MPS representation of the ground state in quantum many-body systems into the RBM wave function consisting of a single multinomial hidden unit. This procedure helped us prepare suitable initial states for the VMC simulation using NNQSs. The computational complexity of the method for converting MPSs into RBM wave functions using the CP decomposition scaled polynomially with the number of variational parameters and sites in the system. The error of the initial wave function typically decreases as $\sim R^{-2}$ with increasing the rank $R$ of the CP decomposition. We examined the applicability of the method by taking the transverse-field Ising model as an example and found that the RBM wave function with a single hidden unit well approximated the ground state of the quantum many-body system. Even when the energy of the initial state was slightly away from that of the true ground state, the VMC method was able to efficiently optimize the parameters in the RBM wave function and provided a better energy estimate after the optimization. We also numerically investigated the required rank $R$ of the CP decomposition for arbitrary system sizes. We empirically found that the condition $R=\mathcal{O}(n)$ is sufficient for any $n$ sites to obtain the ground-state wave function having a small energy error (less than $2\%$ for $R=n/2$) that does not depend on the system size.
Our approach can be extended to more general NNQSs, such as those combining the multinomial RBM [$\Psi_{\mathrm{mRBM}}(\{v_{i}\})$] and conventional binomial RBM [$\Psi_{\mathrm{bRBM}}(\{v_{i}\})$] wave functions:
@@
where $b_{i,j}$, $W_{i,j}$, $a^{\prime}_{i}$, $b^{\prime}_{j}$, and $W^{\prime}_{i,j}$ are variational parameters. The binomial RBM wave function further improves the approximation of the ground state in quantum many-body systems.
In strongly correlated electron systems, the low-energy eigenstates often exhibit complex nodal structures [103]. In most of the VMC simulations, we imitate such nodal structures by using one-body wave functions constructed from the Slater determinant or the Pfaffian, derived from ground states of noninteracting Fermion systems or mean-field solutions of Fermi-Hubbard systems [104, 105]. Our method would provide a complementary way to prepare wave functions with complex nodal structures, which will be more accurate starting points than the one-body wave functions consisting of the Slater determinant or the Pfaffian, once we have accurate tensor networks representing the low-energy eigenstates in strongly correlated electron systems.
@@
Finally, we would like to point out that the CP decomposition of MPSs and tensor trains [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] is a general technique that can be applied to other research fields, such as information science and engineering [107, 108]. One may consider the Monte Carlo approach to obtain the CP decomposition of MPSs or tensor trains using the RBM-type neural network. Indeed, extended Boltzmann machine representations of the CP decomposition have been proposed for nonnegative tensors, formulated as a convex optimization problem based on information geometry recently [109, 110]. Such approach is also an interesting future direction.

### v3

@@ III.2 VMC simulation under open boundary conditions
Figure 11: (a) Energy error as a function of the rank $R$ of the CP decomposition for $h/J=1$ and $n=16$. Filled circles represent the energy error ($|E_{\mathrm{initial,open}}-E_{\mathrm{exact,open}}|/|E_{\mathrm{exact,open}}|$) of the initial state for open boundary conditions used for the VMC method. Up-pointing and down-pointing triangles represent the energy error ($|E_{\mathrm{VMC,open}}-E_{\mathrm{exact,open}}|/|E_{\mathrm{exact,open}}|$, $|E_{\mathrm{VMC,periodic}}-E_{\mathrm{exact,periodic}}|/|E_{\mathrm{exact,periodic}}|$) of the optimized state for open and periodic boundary conditions, respectively. (b) Same as panel (a) but shown on a logarithmic scale. We plot the line proportional to $R^{-2}$ as a reference.
@@ III.3 VMC simulation under periodic boundary conditions
In general, it is hard to prepare the MPS representation of the ground-state wave function under periodic boundary conditions. However, the ground-state wave function should be insensitive to boundary conditions when the system is large enough. Therefore, one may seed the wave function for periodic boundary conditions with the initial wave function for open boundary conditions. We demonstrate that the initial state obtained by the CP decomposition for the open-boundary system can also be efficiently used for the simulation under periodic boundary conditions.
@@ III.4 Size dependence of the required rank of the CP decomposition
Figure 14: Size dependence of the ground-state energy per site for $h/J=0.5$ under periodic boundary conditions. The rank $R$ of the CP decomposition is chosen to be $R=n/2$ for all systems of $n$ sites. Circles and triangles represent the exact and VMC energy per site, respectively. As references, we also plot a dashed line interpolating the exact data points and the line multiplied by $0.999$, corresponding to a $0.1\%$ error, by a dotted line.
Figure 15: Size dependence of the ground-state energy per site for $h/J=2$ under periodic boundary conditions. The rank $R$ of the CP decomposition is chosen to be $R=n/2$ for all systems of $n$ sites. Circles and triangles represent the exact and VMC energy per site, respectively. As references, we also plot a dashed line interpolating the exact data points and the line multiplied by $0.96$, corresponding to a $4\%$ error, by a dotted line.
Finally, we numerically investigate how large the rank $R$ of the CP decomposition should be for arbitrary system sizes. We mainly focus on a rather difficult case at the critical point ($h/J=1$) of the transverse-field Ising model under periodic boundary conditions. The ground state is known to exhibit a logarithmic correction in the entanglement entropy, which naturally requires slightly larger bond dimensions in the MPS representation compared to bond dimensions required for conventional gapped ground states. As we see below, for all systems of $n$ sites, the condition $R=\mathcal{O}(n)$ is sufficient to obtain the ground-state wave function having a small energy error that is independent of the system size.
We first prepare the MPS representation of the ground-state wave function with the maximum bond dimension $D_{\mathrm{max}}=n/2$ for all systems of $n$ sites using the DMRG method. The energy error from the true ground state is found to be less than $10^{-9}$ in the units of $J$ for any system size that we consider. We then CP decompose the MPS representation using the ALS method with $100$ iterations from the random initial CP factors. We optimize the parameters in the RBM wave function generated by the CP decomposed tensor using the VMC method. We keep the rank $R$ of the CP decomposition to $R=D_{\mathrm{max}}(=n/2)$ for all systems of $n$ sites. The number of variational parameters ($b_{l,k}$ and $W_{l,k}$) is $2nR=n^{2}$, which is much smaller than the dimension of the Hilbert space. We use $10^{6}$ ($4\times 10^{6}$) Monte Carlo steps for evaluating physical quantities during the optimization (during the statistical processing).
@@
Away from the critical point ($h/J=1$), the energy error obtained by the VMC method is also independent of the system size $n$ when the condition $R=\mathcal{O}(n)$ is satisfied. We show the ground-state energy per site with $R=n/2$ for $h/J=0.5$ and $h/J=2$ under periodic boundary conditions in Figs. 14 and 15. The VMC method combined with the initial wave function generated by the CP decomposed MPSs performs effectively regardless of whether the system is at the critical point.
In general, at the critical point, the size dependence of the entanglement entropy contains a logarithmic correction, which requires more parameters to accurately represent the corresponding wave function during numerical simulations. Therefore, one may expect the largest energy error at the critical point ($h/J=1$). However, our numerical results suggest that the relative energy error is slightly larger for a parameter region which is away from the critical point. We do not have a clear explanation for the origin of this behavior; however, the RBM wave function does not prevent the efficient representation of quantum states exhibiting entanglement growth that exceeds area-law behavior. Because of the flexible structure, the RBM wave function does not necessarily suffer from the same limitations as the MPS wave function at corresponding points in the parameter space. These results are actually encouraging, because our approach may enable successful simulations in parameter regions that are typically difficult to handle with the MPS wave function.
@@ IV Summary and outlook
We proposed a method for approximately converting the MPS representation of the ground state in quantum many-body systems into the RBM wave function consisting of a single multinomial hidden unit. This procedure helped us prepare suitable initial states for the VMC simulation using NNQSs. The computational complexity of the method for converting MPSs into RBM wave functions using the CP decomposition scaled polynomially with the number of variational parameters and sites in the system. The error of the initial wave function typically decreases as $\sim R^{-2}$ with increasing the rank $R$ of the CP decomposition. We examined the applicability of the method by taking the transverse-field Ising model as an example and found that the RBM wave function with a single hidden unit well approximated the ground state of the quantum many-body system. Even when the energy of the initial state was slightly away from that of the true ground state, the VMC method was able to efficiently optimize the parameters in the RBM wave function and provided a better energy estimate after the optimization, as we confirmed in the case of periodic boundary conditions. This observation suggests that there remains a high likelihood of obtaining an accurate ground state through the VMC method even if the DMRG simulation is inaccurate, causing the corresponding initial state to be slightly away from the ideal state. We also numerically investigated the required rank $R$ of the CP decomposition for arbitrary system sizes. We empirically found that the condition $R=\mathcal{O}(n)$ is sufficient for any $n$ sites to obtain the ground-state wave function having a small energy error (less than $2\%$ for $R=n/2$ at the critical point) that does not depend on the system size.
Our approach can be extended to more general NNQSs, such as those combining the multinomial RBM $\Psi_{\mathrm{mRBM}}(\{v_{i}\})$] and conventional binomial RBM $\Psi_{\mathrm{bRBM}}(\{v_{i}\})$] wave functions:
@@
where $b_{i,j}$, $W_{i,j}$, $a^{\prime}_{i}$, $b^{\prime}_{j}$, and $W^{\prime}_{i,j}$ are variational parameters. The binomial RBM wave function further improves the approximation of the ground state in quantum many-body systems. The conventional binomial RBM wave function would efficiently represent quantum states that exhibit volume-law entanglement entropy [43, 47]. While the multinomial RBM wave function may struggle to capture sufficient entanglement entropy needed for accurately modeling quantum states in some unfortunate cases, combining the binomial RBM wave function can offer a more effective alternative. The integrated RBMs are unlikely to significantly hinder the variational wave function’s convergence toward the ground state when the initial state is already close to the target state.
In strongly correlated electron systems, the low-energy eigenstates often exhibit complex nodal structures [108]. In most of the VMC simulations, we imitate such nodal structures by using one-body wave functions constructed from the Slater determinant or the Pfaffian, derived from ground states of noninteracting fermion systems or mean-field solutions of Fermi-Hubbard systems [109, 110]. The CP decomposition has recently been employed to enhance fermionic trial wave functions within the VMC method [111]. Our method would provide a complementary way to prepare wave functions with complex nodal structures, which will be more accurate starting points than the one-body wave functions consisting of the Slater determinant or the Pfaffian, once we have accurate tensor networks representing the low-energy eigenstates in strongly correlated electron systems.
We have thoroughly examined the performance of the proposed method in the simplest quantum one-dimensional systems to deepen the understanding of how one can efficiently transform tensor network states into shallow RBM wave functions. It is an interesting future direction to extend our method to higher-dimensional systems. In higher dimensions, optimized tensor network states could be inaccurate, causing the corresponding initial RBM wave functions to be far from the ground state. Nevertheless, this initial energy is expected to be much closer to the ground-state energy than those obtained from random initializations and those prepared by the conventional mean-field approximation. Then, the optimization from the initial RBM wave function prepared by our method is more likely to be successful. Investigating the performance of our method in general higher-dimensional systems requires much more careful analyses and will be left for future work.
@@
Note that, even for systems that break the translational symmetry, tensor network states in higher dimensions possess structural and computational characteristics that may differ significantly from the present shallow RBM wave function and the 1D tensor network state. Therefore, generalizing our method to higher dimensions is not straightforward, and one may develop a more specific method for converting PEPSs into RBM wave functions. Moreover, the optimization of PEPS itself remains a challenging and active research field. Addressing these challenges is an interesting future direction.
Finally, we would like to point out that the CP decomposition of MPSs and tensor trains [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] is a general technique that can be applied to other research fields, such as information science and engineering [113, 114]. One may consider the Monte Carlo approach to obtain the CP decomposition of MPSs or tensor trains using the RBM-type neural network. Indeed, extended Boltzmann machine representations of the CP decomposition have been proposed for non-negative tensors, formulated as a convex optimization problem based on information geometry recently [115, 116]. Such approach is also an interesting future direction.
@@ Acknowledgements.
Data availability
The data that support the findings of this article are openly available [117].